OpenStudy (anonymous):

t=2n-2 how to do this pls help

OpenStudy (anonymous):

what are you trying to do here?

OpenStudy (anonymous):

are you solving for a particular variable or what?

OpenStudy (anonymous):

just what to get answer for my maths

OpenStudy (anonymous):

you want to make n the subject?

OpenStudy (anonymous):

ie end up with n = something?

OpenStudy (anonymous):

do u know how to solve this t=2n-2 how to do this

OpenStudy (anonymous):

t=6-3n is it like this answer t=-3n

OpenStudy (anonymous):

thats what u proboably want: t = 2n - 2 add 2 to both sides of the = sign: t + 2 = 2n - 2 + 2 t+ 2 = 2n now divide both sides by 2: (t + 2) /2 = n i put brackets around the t+2 to show that all of it is divided by the 2 answer is n = (t+2 )/ 2

OpenStudy (anonymous):

no t=-3n is not the answer

OpenStudy (anonymous):

i think you are a little confused about what you want all you can do with these is rearrange them to make one variable (in this case t or n) the subject of the expression.

OpenStudy (anonymous):

yes still little confuse i ask my parent to look thx

OpenStudy (anonymous):

can u give me the answer to this one : ) pls

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

t=6-3n

OpenStudy (anonymous):

ok - by a similar method - i only give u the answer as i must go n = (6-t) / 3

OpenStudy (anonymous):

plz click on good answer to close the question - thanx

OpenStudy (anonymous):

ok thx alot bye